# Bilevel optimization with sustainability perspective: a survey on applications

### v2

@@ Abstract
(R) OR in environment and climate change , Bilevel optimization , Mathematical modeling , Sustainability , Applications
@@
[inst1]organization=DEMB, University of Modena and Reggio Emilia, addressline=Via Berengario 51, city=Modena, postcode=41121, country=Italy
@@
[inst2]organization=DISMI, University of Modena and Reggio Emilia, addressline=Via Amendola 2, city=Reggio Emilia, postcode=42122, country=Italy
@@
[inst3]organization=IDS Department, ESSEC Business School, city=Cergy-Pontoise, country=France
@@ 1 Introduction
Sustainability is a broad and multidimensional concept that has evolved significantly over time. A big influence on the use of this term came from the 1983 United Nations Commission’s on Environment and Development Brundtland Report (World Commission 1987) that defined sustainable development as “development that meets the needs of the present without compromising the ability of future generations to meet their own needs”. The United Nations’ Sustainable Development Goals (SDGs) provide now a structured framework for sustainability across economic, environmental, and social dimensions (United Nations 2015). The SDGs are 17 interrelated goals that address a range of global challenges, including poverty eradication, climate action, responsible consumption and production, and sustainable economic growth. These goals build upon the triple bottom line concept (Elkington 2002), which defines sustainability through the three interconnected economic, environmental, and social dimensions.
Based on this principle, economic development should be pursued by companies and societies in a manner that benefits both the well-being of people and the health of the planet. Sustainability-driven decision-making, particularly in policy and industrial settings, requires balancing economic interests with ecological and societal well-being. Bilevel optimization, by explicitly modeling hierarchical decision-making structures, provides a powerful tool for integrating these considerations. By incorporating sustainability principles into optimization models, researchers and decision-makers can explore trade-offs between multiple objectives, addressing several SDGs and facilitating solutions that support sustainable development.
@@
Sustainable bilevel applications Transportation and logistics Green location and transportation Green routing Production planning and manufacturing Carbon policies Green product design and manufacturing Water, waste, and agriculture management Water management Waste management Agriculture Supply chains Green supply chain design and planning Eco-industrial parks Disaster prevention and response Hazmat transportation Hazmat toll setting Hazmat control policies Emergency planning and disaster response
@@
We do not cover single-level optimization or other game theory concepts dealing with sustainability, nor bilevel applications without sustainable applications. We focus on bilevel optimization applications within the operations research literature that apply mathematical programming or known metaheuristic frameworks, and exclude tri-level and multi-level problems, i.e., optimization problems where players are in hierarchical games of three or more levels.
@@ 2 Overview of bilevel applications
where $y$ is the vector of all decisions of all followers and GNEP$(x)$ is the set of generalized Nash equilibra of the non-cooperative game among the $n_{f}$ followers, where the objective function and the feasible set of every follower explicitly depends on the decisions of the other followers. Indicating by $y_{-j}$ the decision variables of all followers except for those of the $j$-th follower, the LL problem of the $j$-th follower can be generally formulated as:
@@
Single-level reformulations. Assuming convexity and a suitable constraint qualification of the LL problem for all possible UL decisions, an SLSF problem is usually reformulated into a single-level problem, that is, a single-level reformulation (SLR) by imposing Karush-Kuhn-Tucker (KKT), strong duality conditions for the follower’s solution $y$, or variational inequalities. In the first case, we will refer to the KKT reformulation (KKTR). Solving the SLR to global optimality guarantees finding an optimistic globally optimal bilevel solution. Another standard reformulation is the lower-level value function reformulation (LLVFR), which relies on the LL optimal value function definition (Outrata 1990). For instance, Fischer et al. (2022) recently proposed a Newton method for finding stationary points of optimistic bilevel problems based on the LLVFR with successful results on well-known instances. Notably, the LLVFR does not require the convexity and constraint qualification assumptions for equivalence between the SLR and the bilevel problem optimality. In general, the value function is not differentiable resulting in a non-smooth optimization problem. The LLVFR guarantees both local and global equivalence to the bilevel problem (Zemkoho and Zhou 2021), whereas KKTR only ensures global equivalence. Indeed, locally optimal solutions of the KKTR are not necessarily locally optimal for the original bilevel problem due to the introduction of implicit variables in the reformulation (Dempe and Dutta 2012). To the best of our knowledge, no definitive preference has been established between KKTR and LLVFR, and both approaches have been applied in the literature. Other duality-based reformulations have been analyzed in Dempe and Mehlitz (2025). There are various techniques for solving such obtained non-convex nonlinear models (i.e., SLRs) to optimality.
@@
Exact methods. Exact solution approaches applied to LP, MILP, and mixed-integer nonlinear programming (MINLP) bilevel problems include Branch-and-Bound (B&B), Branch-and-Cut (B&C), Column Generation (CG), Cutting Plane (CP), or Benders’ Decomposition (BD) algorithms. Typically, a suitable SLR is derived, to which bilevel-specific B&B or B&C techniques are applied. Some duality-based SLRs can also be solved directly using general-purpose solvers.
The penalty function method and approximation heuristics are typically applied to the KKTR. Enumeration algorithms, including the well-known $k$-th best algorithm, are sometimes applied to bilevel linear programs.
For bilevel MILPs, the common approach is to first solve the high-point relaxation and then iteratively discard bilevel infeasible solutions using adapted branching and cutting techniques, see Fischetti et al. (2018); Tahernejad et al. (2020). Benders’ decomposition can be adapted for bilevel MILPs, assuming high-point relaxation compactness (Bolusani and Ralphs 2022). For non-convex bilevel MILPs, a few decomposition and branching approaches have been proposed, see, e.g., Gaar et al. (2024) and further references therein.
@@ 3.1.1 Price setting problems
$\max_{T\in\mathrm{T},y}\quad Ty$ (11)
[…]
@@ 3.1.2 Joint pricing and network design
$\max_{T\in\mathrm{T},\;x\in X,y}\quad Ty$ (13)
[…]
@@ 3.1.3 Pricing with equilibrium constraints
[…]
@@ 3.2   Green location and transportation
Ben-Ayed et al. (1992) and Sinha et al. (2015) introduce transportation bilevel problems focused on highway pricing and network design. These problems fall under the SLSF class, where the government acts as the leader and the average user serves as the follower. Ben-Ayed et al. (1992) present an early bilevel applications in pure network design, addressing rural highway systems in developing countries. The government aims to increase the capacity of the existing highway network, while the follower solves a flow optimization problem. This early application aligns with sustainable development and the SDGs by promoting equitable access to transportation in rural areas, improving infrastructure resilience, and minimizing environmental impact. The authors develop an iterative heuristic algorithm and solve a realistic instance from a case study in Tunisia. Sinha et al. (2015) extend the pricing model (11)-(12c) from Section 3.1.1 for the highway pricing problem with multi-objective UL and LL. In particular, the leader aims to maximize revenues ($F_{1}$, defined in (11)) while minimizing pollution costs ($F_{2}(y,z)=a_{1}y+b_{1}z$). At the LL, the follower aims to minimize total travel costs ($f_{1}$, defined in (12a)) and total travel time ($f_{2}(y,z)=a_{2}y+b_{2}z$). A Pareto-optimal front of decisions balancing revenues and pollution is given for the UL problem, and an LL Pareto-optimal front balancing travel cost and time is given for any UL optimal decision. The authors propose a multi-objective evolutionary algorithm (MOEA) and solve an illustrative example. They also extend the problem to the SLMF version.
@@
In the presence of multiple followers (SLMF), sustainable tolls setting problems are often modeled as bilevel pricing problems with equilibrium (see model (15a)-(15d) of Section 3.1.3). These models are frequently extended to incorporate multiple objectives in the UL, focusing not (only) on the leader’s profit maximization, but on various performance measures of economic, environmental, and health sustainability in transportation. In the UL, the leader is responsible for setting tolls. The followers are the travelers of the system and the LL models an equilibrium with minimum travel time/cost for every traveler. Wang et al. (2014) study the bilevel tolls setting problem with three minimization objectives in the UL (namely, total system travel time, CO2 emissions produced by fuel consumption, and human pollutant uptake), and a bi-objective equilibrium in the LL. The bi-objective equilibrium model is obtained by expressing the objective function $f_{a}(y_{a},T)$ in equation (15a) as the sum of two separable functions $T_{a}(y_{a})$ and $M_{a}(T)$, respectively, for the total travel time and total toll cost, and exploiting results from bi-objective Wardrop equilibrium theory (see Wang and Ehrgott 2013). The authors propose a metaheuristic algorithm where the multi-objective UL problem is addressed by NSGA-II and the LL problem by a quasi-Newton method, and solve an illustrative example, showing the conflicting nature of the three considered UL objectives. In Wen and Eglese (2016), an illustrative example solved by simulating different toll pricing strategies and users’ behavior reveals the impact of various constant and variable pricing strategies on travel time and CO2 emissions. The LL problem solves a dynamic user equilibrium, the simplest dynamic extension of Wardrop’s equilibrium, based on which all routes used by travelers departing at the same time with same origin and destination must have equal and minimal travel time (Szeto and Wong 2012). The authors rely on queuing theory for modeling the LL equilibrium. In the problem proposed by Zhao et al. (2016), the leader aims at setting an optimal GHG emissions charge scheme for travelers (emission tolls), who decide on their travel mode (e.g., cars, motorcycles, bus). The solution method integrates an EA for the UL and the “all-or-nothing” (AON) algorithm (all the demand of each OD pair is assigned to the shortest path for that pair) for the LL traffic assignment. In the multimodal adaptation of LL equilibrium formulation (15a)-(15d), there is an additional objective function term to consider the different travel cost for each travel mode times the related probability of choosing that mode. The authors show managerial insights on the trade-off between the emission reduction goal and system efficiency and the impact of tolls differentiation across travel modes on optimal emission tolls from a numerical example.
@@
In the FLP, the objective is to strategically position a set of facilities to minimize the cost of meeting the demands of a group of customers (see Laporte et al. 2019). In the bilevel FLPs, the leader typically decides on the location and size of the facilities that could be used by the followers. In the cases where the followers are drivers optimizing routes and facility choices, the LL problem is modeled as an equilibrium problem. In green bilevel FLPs, environmental or social decisions and goals are tackled by the leaders. Chen et al. (2020) formulate the bilevel problem of optimal location and capacity planning of electric vehicle (EV) charging stations. The classical LL equilibrium in model (15a)-(15d) is adapted here to consider the total flow on links as the sum of gasoline vehicles and EVs arrivals. In the equilibrium, the travel and waiting time at EV charging stations is minimized for every follower (charging and route choice equilibrium), also considering coupling constraints that restrict flow assignments to only constructed facilities. The authors solve the proposed KKTR with an iterative algorithm (IA), which certificates the C-stationarity of bilevel solutions, on different illustrative examples (up to around 30 nodes, 80 links, and 40 candidate locations). Zhou et al. (2023) consider a bilevel model with a multimodal network equilibrium in the LL for the optimal placement of transit-oriented development (TOD) stations in a Chinese city. TOD is an urban planning approach focused on creating compact and mixed-use urban areas closely linked with mass transit stations to efficiently integrate jobs, housing, and services. The LL is a multimodal stochastic user equilibrium (SUE). The SUE differs from Wardrop equilibrium in the fact that followers make probabilistic mode choices (logit model) to minimize perceived travel times. The authors develop a metaheuristic algorithm and solve a large-size real-world instance. They also show how TOD stations can reduce carbon emissions and how they impact multimodal networks.
@@ 5.1   Water management
Water is a primary source of life and its management has consistently posed complex environmental and societal challenges. Multiple stakeholders are involved in the water management process and different issues may arise. We cover water scarcity, water pollution, and water system network optimization (see Table LABEL:tab:water-man).
@@
In the work of Zhao et al. (2013), an SLMF bilevel programming formulations is employed to model the hierarchical structure of authorities responsible for lake water pollution control. In the UL, there is one leader, the lake authority. In the LL, there are multiple followers, the sub-regional authorities. The leader strategically sets tax rates while ensuring compliance with environmental quality standards. If the pollutant amount generated by a region exceeds the national standard, the region pays a fee proportional to the pollutant quantity. Pollution reduction costs are set by the lake authority. The followers, in turn, make decisions regarding their pollution reduction tactics. The authors derive the KKTR and then apply the bisection method to find an approximated solution to the bilevel problem, providing a global convergence rate.

### v3

@@ Abstract
OR in environment and climate change , Bilevel optimization , Mathematical modeling , Sustainability , Applications
@@
[inst1]organization=DISMI, University of Modena and Reggio Emilia, addressline=Via Amendola 2, city=Reggio Emilia, postcode=42122, country=Italy
@@
[inst2]organization=DEMB, University of Modena and Reggio Emilia, addressline=Via Berengario 51, city=Modena, postcode=41121, country=Italy
@@
[inst3]organization=IDO Department, ESSEC Business School, city=Cergy-Pontoise, country=France
@@ 1 Introduction
Sustainability is a broad and multidimensional concept that has evolved significantly over time. A big influence on the use of this term came from the 1983 United Nations Commission’s on Environment and Development Brundtland Report (World Commission on Environment and Development 1987) that defined sustainable development as “development that meets the needs of the present without compromising the ability of future generations to meet their own needs”. The United Nations’ Sustainable Development Goals (SDGs) provide now a structured framework for sustainability across economic, environmental, and social dimensions (United Nations 2015). The SDGs are 17 interrelated goals that address a range of global challenges, including poverty eradication, climate action, responsible consumption and production, and sustainable economic growth. These goals build upon the triple bottom line concept (Elkington 2002), which defines sustainability through the three interconnected economic, environmental, and social dimensions.
Based on this principle, economic development should be pursued by companies and societies in a manner that benefits both the well-being of people and the health of the planet. Sustainability-driven decision-making, particularly in policy and industrial settings, requires balancing economic interests with ecological and societal well-being. Bilevel optimization, by explicitly modeling hierarchical decision-making structures, provides a powerful tool for integrating these considerations. By incorporating sustainability principles into optimization models, researchers and decision makers can explore trade-offs between multiple objectives, addressing several SDGs, and facilitating solutions that support sustainable development.
@@
Sustainable bilevel applications Transportation and logistics Green location and transportation Sustainable routing Production planning and manufacturing Carbon policies Green product design and manufacturing Water, waste, and agriculture management Water management Waste management Agriculture Supply chains Green supply chain design and planning Eco-industrial parks Disaster prevention and response Hazmat transportation Hazmat toll setting Hazmat control policies Emergency planning and disaster response
@@
We do not cover single-level optimization or other game theory concepts dealing with sustainability, nor bilevel applications without sustainable applications. We focus on bilevel optimization applications within the operations research literature that apply mathematical programming or known metaheuristic frameworks, and exclude multi-level problems, i.e., optimization problems where players are in hierarchical games of three or more levels.
@@ 2 Overview of bilevel applications
where $y$ is the vector of all decisions of all followers and GNEP(x)(x) is the set of generalized Nash equilibra of the non-cooperative game among the $n_{f}$ followers, where the objective function and the feasible set of every follower explicitly depend on the decisions of the other followers. Indicating by $y_{-j}$ the decision variables of all followers except for those of the $j$-th follower, the LL problem of the $j$-th follower can be generally formulated as:
@@
Single-level reformulations. Assuming convexity and a suitable constraint qualification of the LL problem for all possible UL decisions, an SLSF problem is usually reformulated into a single-level problem, that is, a single-level reformulation (SLR) by imposing Karush-Kuhn-Tucker (KKT) conditions, strong duality conditions for the follower’s solution $y$, or variational inequalities. In the first case, we will refer to the KKT reformulation (KKTR). Solving the SLR to global optimality guarantees finding an optimistic globally optimal bilevel solution. Another standard reformulation is the lower-level value function reformulation (LLVFR), which relies on the LL optimal value function definition (Outrata 1990). For instance, Fischer et al. (2022) recently proposed a Newton method for finding stationary points of optimistic bilevel problems based on the LLVFR with successful results on well-known instances. Notably, the LLVFR does not require the convexity and constraint qualification assumptions for equivalence between the SLR and the bilevel problem optimality. In general, the value function is not differentiable resulting in a non-smooth optimization problem. The LLVFR guarantees both local and global equivalence to the bilevel problem (Zemkoho and Zhou 2021), whereas KKTR only ensures global equivalence. Indeed, locally optimal solutions of the KKTR are not necessarily locally optimal for the original bilevel problem due to the introduction of implicit variables in the reformulation (Dempe and Dutta 2012). To the best of our knowledge, no definitive preference has been established between KKTR and LLVFR, and both approaches have been applied in the literature. Other duality-based reformulations have been analyzed in Dempe and Mehlitz (2025). There are various techniques for solving such obtained non-convex nonlinear models (i.e., SLRs) to optimality.
@@
Exact methods. Exact solution approaches applied to LP, MILP, and mixed-integer nonlinear programming (MINLP) bilevel problems include Branch-and-Bound (B&B), Branch-and-Cut (B&C), Column Generation (CG), Cutting Plane (CP), or Benders’ Decomposition (BD) algorithms. Typically, a suitable SLR is derived, to which bilevel-specific B&B or B&C techniques are applied. Some duality-based SLRs can also be solved directly using general-purpose solvers. The penalty function method and approximation heuristics are typically applied to the KKTR. Enumeration algorithms, including the well-known $k$-th best algorithm, are sometimes applied to bilevel linear programs. For bilevel MILPs, the common approach is to first solve the high-point relaxation and then iteratively discard bilevel infeasible solutions using adapted branching and cutting techniques, see Fischetti et al. (2018); Tahernejad et al. (2020). Benders’ decomposition can be adapted for bilevel MILPs, assuming high-point relaxation compactness (Bolusani and Ralphs 2022). For non-convex bilevel MILPs, a few decomposition and branching approaches have been proposed, see, e.g., Gaar et al. (2024) and further references therein.
@@ 3.1.1 Price setting problems
$\hskip-85.35826pt\max_{T\in\mathrm{T},y}\quad Ty$ (11)
$\hskip 56.9055pt\text{s.t.}\quad(y,z)\in\text{arg}\min_{y,z}\quad(c_{1}+T)y+c_{2}z\\$ (12a)
$\hskip 28.45274pt\text{s.t.}\quad g(y,z)\leq 0\\$ (12b)
@@ 3.1.2 Joint pricing and network design
$\hskip-85.35826pt\max_{T\in\mathrm{T},\;x\in X,y}\quad Ty$ (13)
$\hskip 56.9055pt\text{s.t.}\quad(y,z)\in\text{arg}\min_{y,z}\quad C_{1}(y,T)+C_{2}(z)$ (14a)
$\hskip 28.45274pt\text{s.t.}\quad g(y,z)\leq 0$ (14b)
@@ 3.1.3 Pricing with equilibrium constraints
$\min_{y,z}\quad\sum_{a\in A}\int_{0}^{y_{a}}f_{a}(\omega,T)d\omega$ (15a)
$\text{s.t.}\quad\quad y_{a}=\sum_{c\in C}\sum_{\begin{subarray}{c}p\in\mathcal{P}_{c}:\\ a\in p\end{subarray}}z_{p}\quad$ $\forall a\in A$ (15b)
@@ 3.2  Green location and transportation
Ben-Ayed et al. (1992) and Sinha et al. (2015) introduce transportation bilevel problems focused on highway pricing and network design. These problems fall under the SLSF class, where the government acts as the leader and the average user serves as the follower. Ben-Ayed et al. (1992) present early bilevel applications in pure network design, addressing rural highway systems in developing countries. The government aims to increase the capacity of the existing highway network, while the follower solves a flow optimization problem. This early application aligns with sustainable development and the SDGs by promoting equitable access to transportation in rural areas, improving infrastructure resilience, and minimizing environmental impact. The authors develop an iterative heuristic algorithm and solve a realistic instance from a case study in Tunisia. Sinha et al. (2015) extend the pricing model (11)-(12c) from Section 3.1.1 for the highway pricing problem with multi-objective UL and LL. In particular, the leader aims to maximize revenues ($F_{1}$, defined in (11)) while minimizing pollution costs ($F_{2}(y,z)=a_{1}y+b_{1}z$). At the LL, the follower aims to minimize total travel costs ($f_{1}$, defined in (12a)) and total travel time ($f_{2}(y,z)=a_{2}y+b_{2}z$). A Pareto-optimal front of decisions balancing revenues and pollution is given for the UL problem, and an LL Pareto-optimal front balancing travel cost and time is given for any UL optimal decision. The authors propose a multi-objective evolutionary algorithm (MOEA) and solve an illustrative example. They also extend the problem to the SLMF version.
@@
In the presence of multiple followers (SLMF), sustainable tolls setting problems are often modeled as bilevel pricing problems with equilibrium (see model (15a)-(15d) of Section 3.1.3). These models are frequently extended to incorporate multiple objectives in the UL, focusing not (only) on the leader’s profit maximization, but on various performance measures of economic, environmental, and health sustainability in transportation. In the UL, the leader is responsible for setting tolls. The followers are the travelers of the system and the LL models an equilibrium with minimum travel time/cost for every traveler. Wang et al. (2014) study the bilevel tolls setting problem with three minimization objectives in the UL (namely, total system travel time, CO2 emissions produced by fuel consumption, and human pollutant uptake), and a bi-objective equilibrium in the LL. The bi-objective equilibrium model is obtained by expressing the objective function $f_{a}(y_{a},T)$ in equation (15a) as the sum of two separable functions $T_{a}(y_{a})$ and $M_{a}(T)$, respectively, for the total travel time and total toll cost, and exploiting results from bi-objective Wardrop equilibrium theory (see Wang and Ehrgott 2013). The authors propose a metaheuristic algorithm where the multi-objective UL problem is addressed by NSGA-II and the LL problem by a quasi-Newton method, and solve an illustrative example, showing the conflicting nature of the three considered UL objectives. In Wen and Eglese (2016), an illustrative example solved by simulating different toll pricing strategies and users’ behavior reveals the impact of various constant and variable pricing strategies on travel time and CO2 emissions. The LL problem solves a dynamic user equilibrium, the simplest dynamic extension of Wardrop’s equilibrium, based on which all routes used by travelers departing at the same time with same origin and destination must have equal and minimal travel time (Szeto and Wong 2012). The authors rely on queuing theory for modeling the LL equilibrium. In the problem proposed by Zhao et al. (2016), the leader aims at setting an optimal GHG emissions charge scheme for travelers (emission tolls), who decide on their travel mode (e.g., cars, motorcycles, bus). The solution method integrates an EA for the UL and the “all-or-nothing” (AON) algorithm (all the demand of each OD pair is assigned to the shortest path for that pair) for the LL traffic assignment. In the multimodal adaptation of LL equilibrium formulation (15a)-(15d), there is an additional objective function term to consider the different travel cost for each travel mode times the related probability of choosing that mode. The authors provide managerial insights from their numerical example, highlighting the trade-off between emission reduction goals and system efficiency, where travel time costs increase more rapidly with each additional percentage reduction in emissions.
@@
In the FLP, the objective is to strategically position a set of facilities to minimize the cost of meeting the demands of a group of customers (see Laporte et al. 2019). In the bilevel FLPs, the leader typically decides on the location and size of the facilities that could be used by the followers. In the cases where the followers are drivers optimizing routes and facility choices, the LL problem is modeled as an equilibrium problem. In green bilevel FLPs, environmental or social decisions and goals are tackled by the leaders. Chen et al. (2020) formulate the bilevel problem of optimal location and capacity planning of electric vehicle (EV) charging stations. The classical LL equilibrium in model (15a)-(15d) is adapted here to consider the total flow on links as the sum of gasoline vehicles and EVs arrivals. In the equilibrium, the travel and waiting time at EV charging stations is minimized for every follower (charging and route choice equilibrium), also considering coupling constraints that restrict flow assignments to only constructed facilities. The authors solve the proposed KKTR with an iterative algorithm (IA), which certifies the C-stationarity of bilevel solutions, on different illustrative examples (up to around 30 nodes, 80 links, and 40 candidate locations). Zhou et al. (2023) consider a bilevel model with a multimodal network equilibrium in the LL for the optimal placement of transit-oriented development (TOD) stations in a Chinese city. TOD is an urban planning approach focused on creating compact and mixed-use urban areas closely linked with mass transit stations to efficiently integrate jobs, housing, and services. The LL is a multimodal stochastic user equilibrium (SUE). The SUE differs from Wardrop equilibrium in the fact that followers make probabilistic mode choices (logit model) to minimize perceived travel times. The authors develop a metaheuristic algorithm and solve a large-size real-world instance. They also show how TOD stations can reduce carbon emissions and how they impact multimodal networks.
@@ 5.1  Water management
Water is a primary source of life and its management has consistently posed complex environmental and societal challenges. Multiple stakeholders are involved in the water management process and different issues may arise. We cover water scarcity and water pollution (see Table LABEL:tab:water-man).
@@
In the work of Zhao et al. (2013), an SLMF bilevel programming formulation is employed to model the hierarchical structure of authorities responsible for lake water pollution control. In the UL, there is one leader, the lake authority. In the LL, there are multiple followers, the sub-regional authorities. The leader strategically sets tax rates while ensuring compliance with environmental quality standards. If the pollutant amount generated by a region exceeds the national standard, the region pays a fee proportional to the pollutant quantity. Pollution reduction costs are set by the lake authority. The followers, in turn, make decisions regarding their pollution reduction tactics. The authors derive the KKTR and then apply the bisection method to find an approximated solution to the bilevel problem, providing a global convergence rate.